# A Large and Balanced Corpus for Fine-grained Arabic Readability Assessment

### v1

@@ 1 Introduction
Text readability impacts understanding, retention, reading speed, and engagement DuBay (2004). Texts above a student’s readability level can lead to disengagement Klare (1963). Nassiri et al. (2023) highlighted that readability and legibility depend on both external features (e.g., production, fonts) and content. Text leveling in classrooms helps match books to students’ reading levels, promoting independent reading and comprehension Allington et al. (2015). Developing readability models is crucial for improving literacy, language learning, and academic performance. Readability levels have long been a key component of literacy teaching and learning. One of the most widely used systems in English literacy is Fountas and Pinnell Fountas and Pinnell (2006), which employs qualitative measures to classify texts into 27 levels (A to Z+), spanning from kindergarten to adult proficiency. Similarly, Taha-Thomure (2017)’s system for Arabic has 19 levels from Arabic letters \<أ> A to \<ق> Q. These fine-grained levels are designed for pedagogical effectiveness, ensuring young readers experience gradual, measurable progress, particularly in early education (K–6) Barber and Klauda (2020). A key advantage is that they can be easily mapped to coarser levels with fewer categories, which may be more efficient for broader applications in readability research and automated assessments. In this paper we present the Balanced Arabic Readability Evaluation Corpus (Barec) – a large-scale fine-grained readability assessment corpus across a broad space of genres and readability levels. Inspired by the Taha/Arabi21 readability reference Taha-Thomure (2017), which has been instrumental in tagging over 9,000 children’s books, Barec seeks to establish a standardized framework for evaluating sentence-level [^2] We segment paragraphs down to syntactic sentences, and use the term sentence even for small standalone text segments such as phrases and single words (e.g. book titles). Arabic text readability across 19 distinct levels, ranging from kindergarten to postgraduate comprehension. Our contributions are: (a) a large-scale curated corpus with 68K+ sentences (1M+ words) spanning diverse genres; and (b) benchmarking of automatic readability assessment models across multiple granularities, including both fine-grained (19 levels) and collapsed tiered systems (e.g., five-level and three-level scales) to support various research and application needs, aligning with previous Arabic readability frameworks Al Khalil et al. (2018); Al-Khalifa and Al-Ajlan (2010).
@@ 3 Barec Corpus Annotation
In this section, we discuss the guidelines and annotation process. In the next section, we discuss corpus selection and statistics.
@@ 3.1 Barec Guidelines
We present below a summarized version of the Barec annotation guidelines. A detailed account of the adaptation process from Taha-Thomure (2017)’s guidelines is in Anonymous (2024).
@@
The readability level system of Taha-Thomure (2017) uses the Abjad order of Arabic letters for 19 levels: 1-alif, 2-ba, 3-jim, through to 100-qaf. This system emphasizes a finer distinction in the lower levels, where readability is more varied. The Barec pyramid (Figure 1) illustrates the scaffolding of these levels and their mapping to, guidelines components, school grades, and three collapsed versions of level size 7, 5, and 3. All four level types (19-7-5-3) are fully aligned to easy mapping from fine-grained to coarse-grained levels. We present results for these levels in Section 6.
@@
Levels are assessed in six key dimensions. Each of these specify numerous linguistic phenomena that are needed to qualify for being ranked in a harder level. The Cheat Sheet used by the annotators in Arabic and its translation in English are included in Appendix A.
@@
The Barec pyramid (Figure 1) illustrates which aspects are used (broadly) for which levels. For example, Spelling criteria are only used up to level 7-zay, while syntax is used until level 60-sin, and word count is not used beyond level 20-kaf.
@@
Table 2: Summary statistics of the Barec Corpus in sentences.
@@ 3.2 Annotation Team and Process
The annotation process began with A0, who led sentence-level segmentation and initial text flagging and selection. We followed the Arabic sentence segmentation guidelines by Habash et al. (2022a). Subsequently, A1-5 were tasked with assigning readability labels to the individually segmented texts. The annotation was done through a simple Google Sheet interface. A1-5 received folders containing annotation sets, comprising 100 randomly selected sentences each. The average annotation speed was around 2.5 hours per batch (1.5 minutes/sentence). Before starting the annotation, all annotators received rigorous training, including three pilot rounds. These rounds provided opportunities for detailed discussions of the guidelines, helping to identify and address any issues. 16 shared annotation sets (100 sentence each) were included covertly to ensure quality and measure inter-annotator agreement (IAA). Finally, we conducted a thorough second review of the corpus data, resulting in every sentence being checked twice for the first phase (10,658 sentences) before continuing to finish the 68,182K sentences (1M words). In total, the annotators annotated 90K sentences, 24.2% of which is not in the final corpus: 3.6% were deemed problematic (typos and offensive topics); 11.8% were part of the second round of first phase annotation; and 8.9% were part of the IAA efforts, not including their unification. We report on IAA in Section 6.1.
@@ 4 Barec Corpus
In the process of corpus selection, we aimed to cover a wide educational span as well as different domains and topics. We collected the corpus from $1,362$ documents, which we manually categorized into three domains: Arts & Humanities, Social Sciences, and STEM (details in Appendix C.2) and three readership groups: Foundational, Advanced, and Specialized (details in Appendix C.3). Table 2 shows the distribution of the sentences across domains and groups (see Appendix F for document and word level distributions). The distribution across readership levels aligns with the corpus’s educational focus, with a higher-than-usual proportion at foundational levels. Variations across domains reflect differences in the availability of texts and reader interest (more Arts & Humanities, less STEM). The corpus uses documents from $32$ different resources. All selected texts are either out of copyright, within the fair-use limit, or obtained in agreement with publishers. The decision of selecting some of these resources is influenced by the fact that other annotations exist for them. Around 26% of all sentences came from completely new sources that were manually typed to make them digitally usable. All details about the resources are available in Appendix C.
@@
Figure 2 shows the distribution of the three readership groups across all readability levels. As expected, foundational texts strictly dominate the lower levels up to 10-ya, then the presence of advanced and specialized texts starts increasing gradually till the highest level. Specialized texts dominate the highest levels, while the middle levels (20-kaf to 40-mim) include a mix of the three groups with a slight advantage for advanced texts.
@@
We split the corpus into Train ($\simeq$80%), Dev ($\simeq$10%), and Test ($\simeq$10%) at the document level. Sentences from IAA studies are evenly divided between Dev and Test as a special evaluation set as they provide multiple references from different annotators for each example. Also, if other annotations exist for a resource (e.g., CamelTB (Habash et al., 2022b) and ReadMe++ (Naous et al., 2024)), we follow the existing splits. Table 3 shows the corpus splits in the level of documents, sentences, and words. We didn’t end up with perfect 80-10-10 splits due to the IAA and existing corpora exceptions. More details about the splits across readability levels, domains, and readership groups are available in Appendix B.
@@ 5.1 Metrics
Accuracy (Acc)
The percentage of cases where reference and prediction classes match in the 19-level scheme. We addition consider three variants, Acc@7, Acc@5, Acc@3, that respectively collapse the 19-levels into the 7, 5, and 3-level schemes discussed in Section 3.
Adjacent Accuracy (Acc $\pm$1)
@@
An extension of Cohen’s Kappa (Cohen, 1968; Doewes et al., 2023) measuring the agreement between predicted and true labels, but applies a quadratic penalty to larger misclassifications, meaning that predictions farther from the true label are penalized more heavily. We consider Accuracy, Adjacent Accuracy, Average Distance, Quadratic Weighted Kappa as the primary metrics for selecting the best system.
Table 4: An example of a sentence and the corresponding input variants.
@@ 5.4 Loss Functions
Since readability levels exhibit a natural ordering, we explore loss functions that account for the distance between predicted and true labels (Heilman et al., 2008). In addition to standard cross-entropy loss (CE), we experiment with Mean Squared Error (Regression), Ordinal Log Loss (OLL) (Castagnos et al., 2022), Soft Labels Loss (SOFT) (Bertinetto et al., 2020), and Earth Mover’s Distance-based loss (EMD) (L. Hou, 2017), as these have been previously used for ordinal classification tasks. OLL, SOFT, and EMD incorporate a distance matrix $D$ into their formulations to penalize predictions proportionally to their distance from the true label. For simplicity, we define the distance between any two adjacent levels as one, setting $D(i,j)=|i-j|$ for labels i and j. For regression, we round the final output to the nearest readability level to ensure predictions align with the 19 levels.
@@ 5.6 Procedure
Table 5: Inter-Annotator Agreement (IAA) results comparing initial annotations to unified labels. While exact agreement is 72.4% on average, disagreements are within a small range since average ACC±$\pm 1$ is 94.6%.
@@ 6.1 Inter-Annotator Agreement (IAA)
In this section, we report on 14 IAA studies, excluding the three pilots and first two IAAs, which overlapped with annotator training.
@@
The average pairwise exact-match over 19 Barec levels between any two annotators is only 61.4%, which reflects the task’s complexity. Allowing a fuzzy match distance of up to one level raises the match to 74.6%. The overall average pairwise level difference is $0.94$ levels. The average pairwise Quadratic Weighted Kappa 81.3% (substantial agreement) confirms most disagreements are minor Cohen (1968); Doewes et al. (2023).
@@
After each IAA study, the annotators discussed and determined a unified readability level for each sentence in the study. Table 5 presents how each annotator (A1–5) performs against the unified annotations. The results show promising agreement between annotators and the unified annotations. While the average accuracy (Acc) is relatively lower (72.4%), reflecting the difficulty of the task, the other metrics, average Acc $\pm$1 (82.5%), Distance (0.639), and Quadratic Weighted Kappa (88.2%), indicate that differences between individual annotators and the unified labels are minor. Additionally, Acc@7, Acc@5, and Acc@3 results suggest that these differences often occur near the boundaries of the 3, 5, and 7 readability groups, rather than being large deviations.
Table 6: Results comparing different combinations of models and input variants on Barec Dev set. Bold are the best results on each matric.
Table 7: Results comparing different loss function, ensemble methods, and oracle performance on Barec Dev and Test sets. Bold are the best results across individual models and across ensembles.
@@ 6.2 Stage 1 Results
Table 6 presents the results of stage 1, where we evaluate different combinations of pretrained models and input variants using cross-entropy loss. Based on the primary metrics (Acc, Acc $\pm$1, Dist, and QWK), we observe that the AraBERTv02 and AraBERTv2 models generally achieve higher performance across multiple input variants. Among input variants, the Word and D3Tok representations tend to yield better results compared to Lex and D3Lex. Specifically, AraBERTv02 with the Word input achieves the highest Acc (53.6%) and Dist (1.18). Notably, AraBERTv2 is the only model that benefits from the D3Tok input compared to the Word input, showing an improvement across all four primary metrics. We argue that this occurs because AraBERTv2 is the only model in this set that was pretrained on segmented data, making it more compatible with morphologically segmented input. These results suggest that both the choice of input variant and the pretrained model significantly impact performance. Based on a majority vote across the four primary metrics, we select AraBERTv02 with the Word input as the best-performing combinations for stage 2, where we evaluate different loss functions.
6.3 Stage 2 and Ensemble Results
Table 7 presents the results of stage 2, where we evaluate different loss functions using the best-performing pretrained model and input variant combinations from stage 1. To reduce the table size, we only report results for the best loss functions, while the full results are available in Appendix E.1.
Stage 2
Among individual models, the best-performing model is AraBERTv02 with Word input using the OLL15 loss function. However, when considering accuracy-based metrics (Acc, Acc@7, Acc@5, Acc@3), the cross-entropy (CE) model performs better, suggesting that CE is more effective in producing exact matches, whereas OLL15 is better at capturing ordinal relationships.

### v2

@@ 1 Introduction
Text readability impacts understanding, retention, reading speed, and engagement DuBay (2004). Texts above a student’s readability level can lead to disengagement Klare (1963). Nassiri et al. (2023) highlighted that readability and legibility depend on both external features (e.g., production, fonts) and content. Text leveling in classrooms helps match books to students’ reading levels, promoting independent reading and comprehension Allington et al. (2015). Developing readability models is crucial for improving literacy, language learning, and academic performance. Readability levels have long been a key component of literacy teaching and learning. One of the most widely used systems in English literacy is Fountas and Pinnell Fountas and Pinnell (2006), which employs qualitative measures to classify texts into 27 levels (A to Z+), spanning from kindergarten to adult proficiency. Similarly, Taha-Thomure (2017)’s system for Arabic has 19 levels from Arabic letters \<أ> A to \<ق> Q. These fine-grained levels are designed for pedagogical effectiveness, ensuring young readers experience gradual, measurable progress, particularly in early education (K–6) Barber and Klauda (2020). A key advantage is that they can be easily mapped to coarser levels with fewer categories, which may be more efficient for broader applications in readability research and automated assessments. In this paper we present the Balanced Arabic Readability Evaluation Corpus (Barec) – a large-scale fine-grained readability assessment corpus across a broad space of genres and readability levels. Inspired by the Taha/Arabi21 readability reference Taha-Thomure (2017), which has been instrumental in tagging over 9,000 children’s books, Barec seeks to establish a standardized framework for evaluating sentence-level [^3] We use sentence to refer to any standalone text segment, including phrases and single words (e.g., book titles). Arabic text readability across 19 distinct levels, ranging from kindergarten to postgraduate comprehension. Our contributions are: (a) a large-scale curated corpus with 69K+ sentences (1M+ words) spanning diverse genres; and (b) benchmarking of automatic readability assessment models across multiple granularities, including both fine-grained (19 levels) and collapsed tiered systems (e.g., five-level and three-level scales) to support various research and application needs, aligning with previous Arabic readability frameworks Al Khalil et al. (2018); Al-Khalifa and Al-Ajlan (2010).
@@ 3 Barec Corpus Annotation
In this section, we summarize the guidelines and annotation process. For more details, see Habash et al. (2025). In the next section, we discuss corpus selection and statistics.
@@ 3.1 Barec Guidelines
We present below a summarized version of the Barec annotation guidelines. A detailed account of the adaptation process from Taha-Thomure (2017)’s guidelines is in Habash et al. (2025).
@@
The readability level system of Taha-Thomure (2017) uses the Abjad order of Arabic letters for 19 levels: 1-alif, 2-ba, 3-jim, through to 19-qaf. This system emphasizes a finer distinction in the lower levels, where readability is more varied. The Barec pyramid (Figure 1) illustrates the scaffolding of these levels and their mapping to, guidelines components, school grades, and three collapsed versions of level size 7, 5, and 3. All four level types (19-7-5-3) are fully aligned to easy mapping from fine-grained to coarse-grained levels. We present results for these levels in Section 6.
@@
Levels are assessed in six key dimensions. Each of these specify numerous linguistic phenomena that are needed to qualify for being ranked in a harder level. Annotators assign each sentence a readability level based on its most difficult linguistic phenomenon. The Cheat Sheet used by the annotators in Arabic and its translation in English are included in Appendix A.
@@
The Barec pyramid (Figure 1) illustrates which aspects are used (broadly) for which levels. For example, spelling criteria are only used up to level 7-zay, while syntax is used until level 15-sin, and word count is not used beyond level 11-kaf.
@@
Table 2: Summary statistics of the Barec Corpus.
@@ 3.2 Annotation Team and Process
The annotation process began with A0, who led sentence-level segmentation and initial text flagging and selection. We followed the Arabic sentence segmentation guidelines by Habash et al. (2022a). Subsequently, A1-5 were tasked with assigning readability labels to the individually segmented texts. The annotation was done through a simple Google Sheet interface. A1-5 received folders containing annotation sets, comprising 100 randomly selected sentences each. The average annotation speed was around 2.5 hours per batch (1.5 minutes/sentence). Before starting the annotation, all annotators received rigorous training, including three pilot rounds. These rounds provided opportunities for detailed discussions of the guidelines, helping to identify and address any issues. 19 shared annotation sets (100 sentence each) were included covertly to ensure quality and measure inter-annotator agreement (IAA). Finally, we conducted a thorough second review of the corpus data, resulting in every sentence being checked twice for the first phase (10,658 sentences) before continuing to finish the 69,441 sentences (1M words). In total, the annotators annotated 92.6K sentences, 25% of which is not in the final corpus: 3.3% were deemed problematic (typos and offensive topics); 11.5% were part of the second round of first phase annotation; and 10.3% were part of the IAA efforts, not including their unification. We report on IAA in Section 6.1.
@@ 4 Barec Corpus
In the process of corpus selection, we aimed to cover a wide educational span as well as different domains and topics. We collected the corpus from 1,922 documents, which we manually categorized into three domains: Arts & Humanities, Social Sciences, and STEM (details in Appendix C.2) and three readership groups: Foundational, Advanced, and Specialized (details in Appendix C.3). Table 2 shows the distribution of the documents, sentences, and words across domains and groups. The distribution across readership levels aligns with the corpus’s educational focus, with a higher-than-usual proportion at foundational levels. Variations across domains reflect differences in the availability of texts and reader interest (more Arts & Humanities, less STEM). The corpus uses documents from 30 different resources. All selected texts are either out of copyright, within the fair-use limit, or obtained in agreement with publishers. The decision of selecting some of these resources is influenced by the fact that other annotations exist for them. Around 25% of all sentences came from completely new sources that were manually typed to make them digitally usable. All details about the resources are available in Appendix C.
@@
Figure 2 shows the distribution of the three readership groups across all readability levels. As expected, foundational texts strictly dominate the lower levels up to 9-ta, then the presence of advanced and specialized texts starts increasing gradually till the highest level. Specialized texts dominate the highest levels, while the middle levels (10-ya to 14-nun) include a mix of the three groups with a slight advantage for advanced texts.
@@
We split the corpus into Train ($\simeq$80%), Dev ($\simeq$10%), and Test ($\simeq$10%) at the document level. Sentences from IAA studies are divided between all splits. However, We will release the IAA studies as a special set as they provide multiple references from different annotators for each example.LABEL:barec-site Also, if other annotations exist for a resource (e.g., CamelTB (Habash et al., 2022b) and ReadMe++ (Naous et al., 2024)), we follow the existing splits. Table 3 shows the corpus splits in the level of documents, sentences, and words. More details about the splits across readability levels, domains, and readership groups are available in Appendix B.
@@ 5.1 Metrics
Accuracy (Acc19)
The percentage of cases where reference and prediction classes match in the 19-level scheme. We addition consider three variants, Acc7, Acc5, Acc3, that respectively collapse the 19-levels into the 7, 5, and 3-level schemes discussed in Section 3.
Adjacent Accuracy ($\pm$1 Acc19)
@@
An extension of Cohen’s Kappa (Cohen, 1968; Doewes et al., 2023) measuring the agreement between predicted and true labels, but applies a quadratic penalty to larger misclassifications, meaning that predictions farther from the true label are penalized more heavily. We consider Quadratic Weighted Kappa as the primary metric for selecting the best system.
Table 4: Example sentence in different input variants.
@@ 5.4 Loss Functions
Since readability levels exhibit a natural ordering, we explore loss functions that account for the distance between predicted and true labels (Heilman et al., 2008). In addition to standard cross-entropy loss (CE), we experiment with Ordinal Log Loss (OLL) (Castagnos et al., 2022), Soft Labels Loss (SOFT) (Bertinetto et al., 2020), Earth Mover’s Distance-based loss (EMD) (L. Hou, 2017), and Regression using Mean Squared Error (Reg) as these have been previously used for ordinal classification tasks. OLL, SOFT, and EMD incorporate a distance matrix $D$ into their formulations to penalize predictions proportionally to their distance from the true label. For simplicity, we define the distance between any two adjacent levels as one, setting $D(i,j)=|i-j|$ for labels i and j. For regression, we round the final output to the nearest readability level to ensure predictions align with the 19 levels.
@@ 6.1 Inter-Annotator Agreement (IAA)
In this section, we report on 16 IAA studies, excluding the three pilots and first two IAAs, which overlapped with annotator training.
@@
The average pairwise exact-match over 19 Barec levels between any two annotators is only 61.1%, which reflects the task’s complexity. Allowing a fuzzy match distance of up to one level raises the match to 74.4%. The overall average pairwise level difference is 0.94 levels. The average pairwise Quadratic Weighted Kappa 81.8% (substantial agreement) confirms most disagreements are minor Cohen (1968); Doewes et al. (2023).
@@
After each IAA study, the annotators discussed and agreed on a unified readability level for each sentence. On average, the exact match between the annotators and the unified level (Acc19) was 71.7%, reflecting the difficulty of the task. However, the high average $\pm$1 Acc19 (82.3%), low Distance (0.65), and strong Quadratic Weighted Kappa (88.1%) suggest that most disagreements between annotators and the unified labels were minor. For more detailed results on IAA, see Habash et al. (2025).
Input Model Acc19 $\pm$1 Acc19 Dist QWK
Word CamelBERT-msa 54.4% 68.7% 1.20 79.1%
MARBERTv2 53.3% 68.0% 1.20 79.1%
AraBERTv02 55.8% 69.2% 1.17 79.2%
AraBERTv2 51.6% 65.9% 1.32 76.3%
Lex CamelBERT-msa 48.3% 64.4% 1.34 77.1%
MARBERTv2 50.1% 64.9% 1.31 77.0%
AraBERTv02 48.8% 65.4% 1.30 78.5%
AraBERTv2 50.1% 65.4% 1.29 77.7%
D3Tok CamelBERT-msa 54.8% 68.2% 1.21 78.2%
MARBERTv2 54.0% 68.5% 1.20 78.9%
AraBERTv02 54.8% 68.1% 1.22 78.2%
AraBERTv2 56.6% 69.9% 1.14 80.0%
D3Lex CamelBERT-msa 51.1% 65.5% 1.29 78.0%
MARBERTv2 51.6% 65.7% 1.28 78.0%
AraBERTv02 53.3% 68.1% 1.24 78.2%
AraBERTv2 53.2% 67.1% 1.24 78.6%
Table 5: Results comparing different combinations of models and input variants on Barec Dev set. Bold are the best results on each metric.
Loss Acc19 $\pm$1 Acc19 Dist QWK
CE 56.6% 69.9% 1.14 80.0%
EMD 55.3% 70.3% 1.11 81.2%
OLL2 35.2% 70.3% 1.25 82.0%
OLL15 47.3% 71.1% 1.13 82.8%
OLL1 50.8% 71.5% 1.12 81.7%
OLL05 53.1% 68.8% 1.18 79.7%
SOFT2 55.8% 69.8% 1.15 80.0%
SOFT3 56.4% 69.9% 1.14 80.1%
SOFT4 56.4% 69.9% 1.15 79.6%
SOFT5 56.2% 69.5% 1.17 79.3%
Reg 43.1% 73.1% 1.13 84.0%
Table 6: Loss functions comparisons on Barec Dev set. We use AraBERTv2 model and D3Tok input with all loss function. Bold are the best results on each metric.
Input Loss Acc19 $\pm$1 Acc19 Dist QWK Acc7 Acc5 Acc3
Word CE 51.6% 65.9% 1.32 76.3% 61.6% 67.2% 74.0%
Lex CE 50.1% 65.4% 1.29 77.7% 60.6% 66.3% 74.9%
D3Tok CE 56.6% 69.9% 1.14 80.0% 65.9% 70.3% 76.5%
D3Lex CE 53.2% 67.1% 1.24 78.6% 63.6% 69.0% 75.3%
D3Tok EMD 55.3% 70.3% 1.11 81.2% 65.2% 70.0% 76.4%
D3Tok Reg 43.1% 73.1% 1.13 84.0% 61.1% 67.8% 75.9%
Average 46.9% 72.5% 1.11 83.4% 64.0% 70.3% 77.2%
Most Common 56.3% 70.0% 1.13 80.4% 66.3% 70.9% 76.9%
Oracle Combo 75.2% 87.4% 0.50 93.8% 83.2% 85.7% 89.1%
Table 7: Results comparing different loss function, ensemble methods, and oracle performance on Barec Dev set. Bold are the best results across individual models and across ensembles.
Input Loss Acc19 $\pm$1 Acc19 Dist QWK Acc7 Acc5 Acc3
Word CE 51.1% 65.1% 1.31 76.2% 60.7% 65.6% 72.2%
Lex CE 51.2% 66.2% 1.23 78.5% 61.1% 66.2% 74.4%
D3Tok CE 55.9% 70.0% 1.12 80.2% 65.1% 69.4% 75.2%
D3Lex CE 53.7% 67.9% 1.17 79.5% 63.8% 69.1% 74.8%
D3Tok EMD 54.9% 71.4% 1.02 83.7% 64.9% 69.0% 75.2%
D3Tok Reg 41.4% 73.5% 1.11 84.4% 59.4% 65.3% 72.8%
Average 46.0% 73.4% 1.06 84.5% 63.6% 69.4% 75.8%
Most Common 56.2% 70.4% 1.07 81.3% 65.9% 70.0% 75.6%
Oracle Combo 75.9% 87.8% 0.46 94.7% 83.5% 85.7% 88.9%
Table 8: Results comparing different loss function, ensemble methods, and oracle performance on Barec Test set. Bold are the best results across individual models and across ensembles.
@@ 6.2 Stage 1 Results
Table 5 presents the results of stage 1, where we evaluate different combinations of pretrained models and input variants using cross-entropy loss. Based on the all metrics, we observe that the AraBERTv02 and AraBERTv2 models generally achieve higher performance across multiple input variants. Among input variants, the Word and D3Tok representations tend to yield better results compared to Lex and D3Lex. Specifically, AraBERTv2 with the D3Tok input achieves the best scores in all metrics. Notably, AraBERTv2 is the only model that benefits from the D3Tok and D3Lex inputs compared to the Word input, showing an improvement across all metrics. We argue that this occurs because AraBERTv2 is the only model in this set that was pretrained on segmented data, making it more compatible with morphologically segmented input. These results suggest that both the choice of input variant and the pretrained model significantly impact performance. Based on all metrics, we select AraBERTv2 with the D3Tok input as the best-performing combination. In stage 2, we evaluate it with different loss functions. The confusion matrix for this model is available in the Appendix D.1.
6.3 Stage 2 Results
Table 6 presents the results of stage 2, where we use the best model from stage 1 to evaluate different loss functions. Among all the loss functions evaluated, Cross-Entropy (CE) achieves the highest exact accuracy (Acc19) at 56.6%, indicating that it performs best when predicting the exact readability level. In contrast, other loss functions show stronger performance on metrics that consider the ordinal nature of readability levels. Notably, Regression achieves the highest $\pm$1 accuracy at 73.1% and the best Quadratic Weighted Kappa (QWK) at 84.0%, suggesting it excels at predicting levels close to the gold label, despite being the worst in terms of exact accuracy. These findings support that loss functions designed for ordinal or continuous labels—such as EMD, OLL, and Regression—are more effective on evaluation metrics that reward proximity to the correct label, even if they underperform on strict accuracy. More results for other loss functions are in Appendix D.2.
6.4 Ensemble Results
Table 7 presents results from Stage 1, where AraBERTv2 is evaluated with four different input variants, and Stage 2, where it is trained using the two best-performing loss functions. It also includes results from two ensemble strategies applied across all six models to assess whether combining predictions can further improve performance. We also include an oracle combination, which represents an upper bound on performance. This allows us to estimate the maximum potential gain achievable through ensembling.